# EGG-SR: Embedding Symbolic Equivalence into Symbolic Regression via Equality Graph

### v1

@@ 3.2 Egg-SR: Embed Symbolic Equivalence into Symbolic Regression via Egg
Figure 2: Execution pipeline of our Egg-MCTS. (a) Starting at the root of the tree, the algorithm selects the child with the highest UCT score (in equation 2) until reaching a leaf. (b) The selected leaf is expanded by applying all applicable production rules, producing new child nodes. (c) For each child, several simulations are run to complete the expression template by sampling additional rules. The resulting expressions are fitted to the data to estimate their coefficients. (d) Rewards and visit counts from evaluated children are back-propagated up the tree. Updates are applied to the selected and also equivalent paths (highlighted in two colors), enabled by our Egg module.
@@
Embedding Egg into Monte Carlo Tree Search. MCTS (DBLP:conf/iclr/Sun0W023) maintains a search tree to explore optimal expressions. Each edge is labeled with a production rule, and the label of a node is determined by the sequence of edge labels from the root. By construction, the node label corresponds to either a valid or a partially complete expression. A node representing a partially complete expression can be further expanded by extending its children using production rules.
@@
During learning, the MCTS algorithm executes the following four steps (DBLP:journals/kbs/BrenceTD21): (1) Selection. Starting from the root node, successively select the production rule with the highest Upper Confidence Bound for Trees (UCT) (DBLP:conf/ecml/KocsisS06) to reach a leaf node $s$:
$\mathtt{UCT}(s,a)=\mathtt{Reward}(s,a)+c\sqrt{{\log(\mathtt{visits}(s))}/{\mathtt{visits}(s,a)}}$ (2)
@@
where $\mathtt{Reward}(s,a)$ is the average reward obtained by applying production rule $a$ at node $s$, $\mathtt{visits}(s)$ is the total visits to node $s$, and $\mathtt{visits}(s,a)$ is the number of times rule $a$ has been selected at $s$. The constant $c$ is used to balance exploration and exploitation. (2) Expansion. At the selected leaf node $s$, expand the search tree by applying all applicable production rules to generate its child nodes. (3) Simulation. For each child of $s$, conduct a fixed number of rollouts. In each rollout, generate a valid expression $\phi$ by randomly applying production rules. Then, estimate the optimal coefficients in $\phi$ and evaluate its fitness on the training data $D$. The reward for this rollout is typically defined as ${1}/{(1+\mathtt{NMSE}(\phi))}$. (4) Backpropagation. Update the rewards and visit counts for node $s$ and all its ancestors up to the root node. After the final iteration, the expression with the highest fitness score, encountered during rollouts, is returned as the best expression.
Egg-based Backpropagation. At the backpropagation step, we obtain the path from the root to the selected node $s$ along with the estimated reward. Using Egg, we identify equivalent paths and update their estimated rewards simultaneously. This procedure effectively propagates information as if rollouts had been performed on all equivalent subtrees. Our approach is related to prior work on equivalence-aware search in MCTS (DBLP:conf/acml/LeurentM20; DBLP:conf/aips/CzechKK21). Concretely, the path—representing a partially completed expression—is first converted into an initial e-graph. The e-graph is then saturated by repeatedly applying the set of rewrite rules. From this saturated e-graph, we sample several distinct equivalent sequences and check if the tree contains corresponding paths. If so, we apply backpropagation to all of them. In this way, equivalent expressions share the same reward information, used to avoid redundant search. Theorem 3.1 shows that Egg-MCTS accelerates learning by reducing the effective branching factor compared with standard MCTS, meaning fewer child nodes are explored at each learning iteration.
@@
An example execution pipeline of Egg-MCTS is provided in Figure 2. In Figure 2(d), we examine two distinct paths within the search tree maintained by the algorithm:
$\text{Path }1:$ $\;(A\to\log(A),A\to A\times A,A\to x_{1}),\!$ $\text{Node }1:$ $\;s_{1}=\log(x_{1}\times A).$
$\text{Path }2:$ $\;(A\to A+A,A\to\log(A),A\to x_{1},A\to\log(A)),\!$ $\text{Node }2:$ $\;s_{2}=\log(x_{1})+\log(A).$
@@
Here, each path is a sequence of edge labels from the root to the leaf node $s_{i}$. Based on the grammar definition in section 2, node $s_{1}$ corresponds to the partially completed expression $\log(x_{1}\times A)$, while node $s_{2}$ represents $\log(x_{1})+\log(A)$. The two nodes are equivalent under the rewrite rule $\log(ab)\leadsto\log a+\log b$. Consequently, their rewards, estimating the averaged goodness-of-fit of expressions on the dataset, should be approximately equal:
@@
$\mathtt{Reward}(s_{1},a)\approx\mathtt{Reward}(s_{2},a),\qquad\forall\text{ rule }a\in\text{ the set of available rules}$
Exploring sub-trees of $s_{1}$ and $s_{2}$ independently leads to redundant computation, as MCTS repeatedly traverses equivalent regions of the search space, slowing down learning. With Egg-MCTS, the visit counts and reward estimates of both paths are updated simultaneously, eliminating the need for additional rollouts on the green leaf in Figure 2(d) in future rounds.
@@
Embedding Egg into Deep Reinforcement Learning. DRL employs a neural network-based sequential decoder to predict a sequence of production rules, which are then converted into valid expressions according to the grammar. The reward function assigns high values to predicted expressions that fit the training data well. The model is trained to maximize the expected reward of generated expressions evaluated on the training data (DBLP:journals/corr/abs-1801-03526; DBLP:conf/iclr/PetersenLMSKK21; landajuelaunified; ijcai2024p651).
@@
At each step, a neural network-based sequential decoder samples the next production rule from a probability distribution over all available rules, conditioned on the previously generated sequence. The model emits a sequence $\tau$ with probability $p_{\theta}(\tau)$. The reward function is typically defined as $\mathtt{reward}(\tau)={1}/{(1+\mathtt{NMSE}(\phi))}$. The learning objective is to maximize the expected reward: $J(\theta):=\mathbb{E}_{\tau\sim p_{\theta}}\left[\mathtt{reward}(\tau)\right]$, whose gradient is $\mathbb{E}_{\tau\sim p_{\theta}}\left[\mathtt{reward}(\tau)\nabla_{\theta}\log p_{\theta}(\tau)\right]$. In practice, the expectation is approximated via Monte Carlo. Drawing $N$ sequences $\{\tau_{1},\ldots,\tau_{N}\}$ from the decoder, we obtain the policy gradient estimator:
$g(\theta)\approx\frac{1}{N}\sum\nolimits_{i=1}^{N}(\mathtt{reward}(\tau_{i})-b)\nabla_{\theta}\log p_{\theta}(\tau_{i}),$ (3)
@@
where $b$ is a baseline to reduce estimator variance (DBLP:conf/uai/WeaverT01). Recently, (DBLP:conf/iclr/PetersenLMSKK21) proposes to use the Top-$25\%$ quantile of samples to replace the classic empirical mean.
@@
Egg-based Policy Gradient Estimator. For each sampled sequence $\tau_{i}$, we construct an e-graph that compactly encodes all of its equivalent expressions. From this e-graph, we extract K1K-1 equivalent sequences $\{\tau_{i}^{(2)},\ldots,\tau_{i}^{(K)}\}$. The estimator is then revised to
@@
$g_{\mathtt{egg}}(\theta)\approx\frac{1}{N}\sum\nolimits_{i=1}^{N}\big(\mathtt{reward}(\tau_{i})-b^{\prime}\big)\nabla_{\theta}\log\left[\sum\nolimits_{k=1}^{K}p_{\theta}(\tau_{i}^{(k)})\right],$ (4)
@@
where $\tau_{i}^{(1)}$ is the original sequence $\tau_{i}$ and $b^{\prime}$ is the corresponding baseline, and $\sum_{k=1}^{K}p_{\theta}(\tau_{i}^{(k)})$ aggregates the probabilities of all equivalent sequences that share the same reward. The pipeline of DRL and Egg-DRL is presented in Figure 7. Theorem 3.2 shows that Egg-DRL achieves a lower-variance estimator compared to the estimator in standard DRL (DBLP:conf/iclr/PetersenLMSKK21).
@@
Embed Egg into Large-Language Model. LLM is applied to search for optimal symbolic expressions with prompt tuning (DBLP:journals/iclr/llmsr25; DBLP:conf/acl/MerlerHDM24). The procedure consists of three key steps: (1) Hypothesis Generation: The LLM generates multiple candidate expressions based on a prompt describing the problem background and the definitions of each variable. (2) Data-Driven Evaluation: Each candidate expression is evaluated based on its fitness on the training dataset. (3) Experience Management: In subsequent iterations, the LLM receives feedback in the form of previously predicted expressions and their corresponding fitness scores, allowing it to refine future generations. High-fitness expressions are retained and updated over multiple rounds of iteration.
@@
Egg-based Feedback Prompt. Since LLMs typically generate Python functions rather than symbolic expressions directly, we introduce a wrapper that parses the generated Python code into symbolic expressions. These expressions are then transformed into e-graphs using a set of rewrite rules. From each e-graph, we extract semantically equivalent expressions and summarize them into a concise message, which is incorporated into the prompt for the next round. This augmentation enables the LLM to observe a richer set of functionally equivalent expressions, potentially improving the quality and accuracy of subsequent predictions.
3.3 Theoretical Insight on Egg-SR Accelerating Learning
@@
Theorem 3.1 establishes that augmenting MCTS with Egg yields an asymptotically tighter regret bound than standard MCTS, as the effective branching factor satisfies $\kappa_{\infty}\leq\kappa$. Also, Theorem 3.2 shows that embedding Egg into DRL produces an unbiased gradient estimator while strictly reducing gradient variance, ensuring more stable and efficient policy updates.
@@
Consider the MCTS learning framework augmented with Egg. As defined in Definitions 1 and 3, let $T$ denote the total number of learning iterations, $\gamma\in(0,1)$ the discount factor of the corresponding Markov decision process, $\kappa$ be the near-optimal branching factor of standard MCTS, and $\kappa_{\infty}$ the corresponding branching factor of Egg-MCTS. Then the regret bounds satisfy
[…]
@@
Our proof builds on part of the analysis of (DBLP:conf/acml/LeurentM20, Theorem 16), which employs a graph structure to replace the search tree to aggregate identical states. Their proof requires unrolling the graph into a tree that contains every path traversable in the graph. The search tree in our Egg-MCTS behaves almost identically to this unrolled tree. Following their regret analysis of the unrolled tree, we obtain the final regret bounds. A full proof is provided in Appendix B. ∎
@@
Consider embedding Egg into the DRL framework. Let $N$ samples be drawn from the distribution $p_{\theta}$, and Egg module provides additional K1K-1 equivalent variants for each sample. (1) Unbiasedness. The expectation of the standard estimator $g(\theta)$ equals that of the Egg-based estimator $g_{\mathtt{egg}}(\theta)$: $\mathbb{E}_{\tau\sim p_{\theta}}\big[g(\theta)\big]=\mathbb{E}_{\tau\sim p_{\theta}}\big[g_{\mathtt{egg}}(\theta)\big]$. (2) Variance Reduction. The variance of the proposed estimator is smaller than that of the standard estimator: $\mathrm{Var}_{\tau\sim p_{\theta}}\big[g_{\mathtt{egg}}(\theta)\big]\leq\mathrm{Var}_{\tau\sim p_{\theta}}\big[g(\theta)\big]$.
@@
To show unbiasedness, we follow the definition of the estimators to verify that the expected value of $g_{\mathtt{egg}}(\theta)$ equals that of $g(\theta)$. The key idea behind the variance reduction is that Egg groups together equivalent sequences that share the same reward. Grouping sequences with identical rewards leads to lower within-group variance. Full proof is provided in Appendix C. ∎

### v2

@@ 3.2 Embedding Symbolic Equivalence into Symbolic Regression via Egg
Figure 2: The Egg-MCTS pipeline consists of: (a) Starting at the root node, MCTS selects the child with the highest UCT score (in equation 2) until reaching a leaf. (b) The selected leaf produces new child nodes by applying all applicable production rules. (c) For each child, run several rollouts to complete the expression template by sampling additional rules. The resulting expressions are fitted to the data to estimate their coefficients. (d) Rewards and visit counts from the selected leaf are back-propagated to the root. In addition to updating the selected path (blue), we also update those equivalent paths (orange) identified by our Egg module.
@@
Embedding Egg into Monte Carlo Tree Search. MCTS [DBLP:conf/iclr/Sun0W023, ruan2025discovering] maintains a search tree to explore an optimal sequence of decisions, here corresponding to a sequence of production rules. Each edge is labeled with a production rule, and each node is labeled by the sequence of edge labels from the root. By the grammar definition, this node label corresponds to a partially completed (or complete) expression.
@@
During learning, MCTS iterates the following four steps [DBLP:journals/kbs/BrenceTD21]: (1) Selection. Starting from the root node, successively select the edge of node $s$ (noted as $a$) with the highest Upper Confidence Bound for Trees (UCT) [DBLP:conf/ecml/KocsisS06]:
$\mathtt{UCT}(s,a)=\mathtt{reward}(s,a)+\alpha\sqrt{{\log(\mathtt{visits}(s))}/{\mathtt{visits}(s,a)}}$ (2)
@@
where $\mathtt{reward}(s,a)$ is the average reward obtained by selecting edge $a$ at node $s$, $\mathtt{visits}(s)$ is the number of visits to node $s$, and $\mathtt{visits}(s,a)$ is the number of times rule $a$ has been selected at node $s$. The constant $\alpha$ (often set to $\sqrt{2}$ in theory) balances between exploration and exploitation. (2) Expansion. When reaching a leaf node $s$, expand the search tree by generating its child nodes using all production rules. (3) Simulation. Perform several rollouts for each child to evaluate the average reward of node $s$. In each rollout, generate a valid expression $\phi$ by randomly applying production rules until completion. Then, estimate the optimal coefficients in $\phi$ and evaluate its reward. A common reward function is ${1}/{(1+\mathtt{NMSE}(\phi))}$. (4) Backpropagation. Update the reward estimates and visit counts for node $s$ and all its parents up to the root node. After the final iteration, MCTS returns the expression with the highest reward encountered during training as its prediction.
Egg-based Backpropagation.Our backpropagation strategy is motivated by transposition tables [DBLP:conf/cig/ChildsBK08, DBLP:conf/acml/LeurentM20], which use a table to cache identical nodes (e.g., via hashing) in a search tree and share their statistics during training. This mechanism propagates information as if the search had visited all identical nodes. Such tables are effective in domains such as Go and Hearthstone, where nodes that are identical can be easily determined. In symbolic regression, however, two nodes may be identical only up to symbolic equivalence induced by rewrite rules, so a hashing-based transposition table is not directly applicable. To address this, Egg is used to identify equivalent paths and nodes.
Concretely, the path—representing a partially completed expression—is first converted into an initial e-graph. The e-graph is then saturated by repeatedly applying the set of rewrite rules. From this saturated e-graph, we sample several distinct equivalent sequences and check if the tree contains corresponding paths. If so, we apply backpropagation to all of them. In this way, we avoid redundant exploration by sharing the rewards and visit counts of equivalent paths and nodes.
This modification in Egg-MCTS changes the interpretation of equation (2). $\mathtt{visits}(s)$ no longer counts how many times the specific tree node $s$ appears on simulated paths; instead, it counts visits to any representative within the associated equivalence class. Conceptually, this mirrors the transposition table that shares statistics across identical tree nodes.
In Theorem 3.1, we show that Egg-MCTS accelerates learning by reducing the search tree’s effective branching factor relative to standard MCTS. It prevents redundant exploration of equivalent subtrees and concentrates sampling on genuinely distinct (and potentially near-optimal) paths.
Example 3.2.
@@
Figure 2 shows an example execution pipeline of Egg-MCTS. Specifically, Figure 2(d) highlights two distinct root-to-node paths in the search tree:
$\text{Path }1:$ $\;(A\to\log(A),A\to A\times A,A\to x_{1}),$ Node $s_{1}$: $\log(x_{1}\times A).$
$\text{Path }2:$ $\;(A\to A+A,A\to\log(A),A\to x_{1},A\to\log(A)),$ Node $s_{2}$: $\log(x_{1})+\log(A).$
@@
Here, each path is a sequence of edge labels from the root to the leaf node $s_{i}$. Based on the grammar definition in section 2, node $s_{1}$ corresponds to the partially completed expression $\log(x_{1}\times A)$, while node $s_{2}$ represents $\log(x_{1})+\log(A)$. The two nodes are equivalent under the rewrite rule $\log(ab)\leadsto\log a+\log b$. Consequently, their rewards, estimating the averaged goodness-of-fit of expressions on the training data, should be approximately equal:
@@
$\mathtt{reward}(s_{1},a)\approx\mathtt{reward}(s_{2},a),\qquad\forall a\in\text{ the set of production rules}$
Standard MCTS would explore the subtrees rooted at $s_{1}$ and $s_{2}$ independently, because it is unaware of their equivalence. This results in redundant computation and slows down learning. With Egg-MCTS, the visit counts and reward estimates of both paths are updated simultaneously, eliminating the need for extra iterations on the orange leaf in Figure 2(d). See Example B.1 for the case of other rewrite rules, e.g., $\sin^{2}(a)+\cos^{2}(a)\leadsto 1$ and $a/a\leadsto 1$.
@@
Embedding Egg into Deep Reinforcement Learning. DRL typically employs a neural sequential decoder to predict an expression by sampling a sequence of production rules from the model distribution. The reward assigns higher values to expressions that better fit the training data [DBLP:conf/iclr/PetersenLMSKK21, landajuelaunified, ijcai2024p651]. The pipeline of DRL and Egg-DRL is presented in Figure 8.
@@
At every step, the sequential decoder samples the next production rule from a distribution over all available rules, conditioned on the previously generated sequence. The decoder thus induces a distribution $p_{\theta}(\tau)$ over the rule sequence $\tau$. The reward function is typically defined as $\mathtt{reward}(\tau)={1}/{(1+\mathtt{NMSE}(\phi))}$, where $\phi$ is the expression constructed by $\tau$ following grammar definition (in section 2). The learning objective is to maximize the expected reward of generated expressions on the training data: $\mathbb{E}_{\tau\sim p_{\theta}}\left[\mathtt{reward}(\tau)\right]$, whose gradient is $\mathbb{E}_{\tau\sim p_{\theta}}\left[\mathtt{reward}(\tau)\nabla_{\theta}\log p_{\theta}(\tau)\right]$. In practice, the gradient is approximated via Monte Carlo. Sampling $N$ sequences $\{\tau_{1},\ldots,\tau_{N}\}$ from the decoder, the policy gradient estimator computes:
$g(\theta)\approx\frac{1}{N}\sum_{i=1}^{N}(\mathtt{reward}(\tau_{i})-b)\nabla_{\theta}\log p_{\theta}(\tau_{i}),$ (3)
@@
where $b$ is a baseline used to reduce variance [DBLP:conf/uai/WeaverT01]. Recent work [DBLP:conf/iclr/PetersenLMSKK21] further proposes using a top-quantile subset of samples, rather than the sample mean.
@@
Egg-based Policy Gradient Estimator. For each sampled sequence $\tau_{i}$, we construct an e-graph that compactly encodes all of its equivalent expressions. From this e-graph, we sample K1K-1 equivalent sequences $\{\tau_{i}^{(2)},\ldots,\tau_{i}^{(K)}\}$. We then revise the policy-gradient estimator as
@@
$g_{\mathtt{egg}}(\theta)\approx\frac{1}{N}\sum_{i=1}^{N}\big(\mathtt{reward}(\tau_{i})-b^{\prime}\big)\nabla_{\theta}\log\left[\sum_{k=1}^{K}p_{\theta}(\tau_{i}^{(k)})\right],$ (4)
@@
where $\tau_{i}^{(1)}$ is the original sequence $\tau_{i}$ and $b^{\prime}$ is the corresponding baseline, and $\sum_{k=1}^{K}p_{\theta}(\tau_{i}^{(k)})$ aggregates the probabilities of all equivalent sequences that share the same reward. In Theorem 3.2, we show that Egg improves DRL training by yielding a lower-variance gradient estimator than standard DRL [DBLP:conf/iclr/PetersenLMSKK21].
@@
Embed Egg into Large-Language Model. LLM is applied to search for optimal symbolic expressions with prompt tuning [DBLP:conf/acl/MerlerHDM24, DBLP:journals/iclr/llmsr25]. The procedure consists of three key steps: (1) Hypothesis Generation: The LLM generates multiple candidate expressions based on a prompt describing the problem background and the definitions of each variable. (2) Data-Driven Evaluation: Each candidate expression is evaluated based on its fitness on the training dataset. (3) Experience Management: In subsequent iterations, the LLM receives feedback in the form of previously predicted expressions and their corresponding fitness scores, allowing it to refine future generations. High-fitness expressions are retained and updated over multiple rounds of iteration.
@@
Egg-based Feedback Prompt. Since LLMs typically generate Python functions rather than symbolic expressions directly, we introduce a wrapper that parses the generated Python code into symbolic expressions. These expressions are then transformed into e-graphs using a set of rewrite rules. From each e-graph, we extract semantically equivalent expressions and summarize them into a similar feedback message, which is incorporated into the prompt for the next round. This augmentation enables the LLM to observe a richer set of functionally equivalent expressions, potentially improving the quality and accuracy of predictions in future iterations.
3.3 Connection to Existing Methods
Prior work has explored alternative expression representations based on layer-wise symbolic networks (SymNet) [DBLP:conf/icml/SahooLM18, DBLP:conf/icml/LiLYWSLLWDH24], which are not directly compatible with our grammar-based formulation. Recent studies on SymNet further show that many learned coefficients can be aggregated or merged [DBLP:journals/corr/abs-2401-15103]. Extending this notion of coefficient equivalence to sub-expression equivalence within SymNet remains an interesting open problem.
For DRL-based approaches, several extensions of the original method [DBLP:conf/iclr/PetersenLMSKK21] have been proposed, including [mundhenk2021seeding, landajuelaunified]. An important open question is whether symbolic-equivalence can be integrated into these extensions in a compatible and effective manner, and whether doing so can further improve overall performance.
Finally, DBLP:conf/nips/KamiennydLC22, DBLP:conf/nips/ShojaeeMFR23 encode data directly with a Transformer and predict an expression traversal sequence end-to-end under a cross-entropy objective. A natural way to incorporate Egg is to use it during training to generate multiple equivalent, correct target sequences. How to best leverage Egg at inference time, however, remains an open problem.
3.4 Theoretical Justification on Egg-SR Accelerating Learning
Theorem 3.1 shows that Egg-MCTS achieves an asymptotically tighter regret bound than standard MCTS, as the effective branching factor satisfies $\kappa_{\infty}\leq\kappa$. Intuitively, by identifying symbolically equivalent nodes and sharing their search statistics, Egg prevents redundant exploration of equivalent subtrees and concentrates sampling on genuinely distinct (and potentially near-optimal) paths. After many iterations, Egg-MCTS concentrates more quickly on the near-optimal region of the search space, which is captured by a smaller effective branching factor.
@@ 3.4 Theoretical Justification on Egg-SR Accelerating Learning
Also, Theorem 3.2 shows that embedding Egg into DRL produces an unbiased gradient estimator while strictly reducing gradient variance, ensuring more stable and efficient policy updates.
@@
Consider embedding Egg into the MCTS framework. Given Definitions 1 and 3 (in appendix), let $n$ denote the total number of learning iterations, $\gamma\in(0,1)$ the discount factor of the corresponding Markov decision process, $\kappa$ be the near-optimal branching factor of standard MCTS, and $\kappa_{\infty}$ the corresponding branching factor of Egg-MCTS. Then the regret bounds satisfy
$\mathtt{regret}(n)\;=\;\widetilde{\mathcal{O}}\left(n^{-\frac{\log(1/\gamma)}{\log\kappa}}\right),\qquad\mathtt{regret}_{\mathtt{egg}}(n)\;=\;\widetilde{\mathcal{O}}\left(n^{-\frac{\log(1/\gamma)}{\log\kappa_{\infty}}}\right),\qquad\text{with }\kappa_{\infty}\leq\kappa.$
@@
DBLP:conf/acml/LeurentM20 analyze MCTS on a graph obtained by merging identical tree nodes and sharing their statistics. Their analysis unrolls the graph into a tree that contains all graph-traversable paths. The search tree in Egg-MCTS behaves identically to the unrolled tree. Our final results follow their regret analysis on the unrolled tree. A detailed proof is in Appendix B.2. ∎
@@
Consider embedding Egg into the DRL framework. Let $\tau\sim p_{\theta}$ denote trajectories sampled from the distribution $p_{\theta}$, and consider the two estimators defined in Equations (3) and (4). (1) Unbiasedness. The expectation of the standard estimator $g(\theta)$ equals that of the Egg-based estimator $g_{\mathtt{egg}}(\theta)$: $\mathbb{E}_{\tau\sim p_{\theta}}\big[g(\theta)\big]=\mathbb{E}_{\tau\sim p_{\theta}}\big[g_{\mathtt{egg}}(\theta)\big]$. (2) Variance Reduction. The variance of the proposed estimator is smaller than that of the standard estimator:
$\mathbb{V}\mathrm{ar}_{\tau\sim p_{\theta}}\big[g_{\mathtt{egg}}(\theta)\big]\leq\mathbb{V}\mathrm{ar}_{\tau\sim p_{\theta}}\big[g(\theta)\big].$
@@
For (1), unbiasedness can be obtained by expanding the definitions of $g(\theta)$ and $g_{\mathtt{egg}}(\theta)$. For (2), the key observation is that Egg groups together equivalent trajectories that share the same reward. Averaging over sequences with identical rewards reduces within-group variability, which yields a smaller variance. A full proof is provided in Appendix B.3. ∎